Question: 2 following FX rates & interest rates   

CURRENCY SPOT RATES

INTEREST RATES 3M (84 DAYS) (%)

INTEREST RATES 6M (181 DAYS) (%)

EURO/USD

1.0607/15

EUR 3 MONTH

0.02/0.04

EUR 6M

0.1/0.15

GBP/USD

1.5089/106

GBP 3 MONTH

0.25/0.30

GBP 6M

0.30/0.35

AUD/USD

0.7225/33

AUD 3 MONTH

1.05/10

AUD 6M

1.25/35

USD/JPY

122.55/72

JPY 3 MONTH

0.10/15

JPY 6M

0.15/20

USD/CNY

6.3920/45

CNY 3 MONTH

3.80/90

CNY 6M

4.10/20

USD/PKR

106.10/20

PKR 3 MONTH

5.90/00

PKR 6M

6.10/20

USD 3 MONTH

0.20/25

USD 6M

0.40/45

Calculate the Following:

  1. 3 month forward rate & swap points of Euro/AUD
  2. 3 Month forward points and 6 month forward points of JPY/PKR
Answer #1

Answer

(1)

3 months forward rate for Euro/AUD

Spot Rate for Euro/USD is 1.0607/15, which is equivalent to 1.0607/1.0615

Spot Rate for AUD/USD is 0.7225/33, which is equivalent to .07225/0.7233

The 3 months rate for Euro is at a premium 0.02/0.04 and therefore the 3 months forward rate for Euro/USD is calculated as shown below:

Spot: 1.0607/1.0615

add premium: .02/.04

thus, 3 months forward rate is 1.0807/1.1015 for Euro /USD.

Similarly, the 3 months rate for AUD is at a premium 1.05/1.10 and hence the 3 months forward rate for AUD/USD is calculated as shown below:

spot = 0.7225/0.7233

add premium:=1.05/1.10

thus, months forward rate is 1.7725/1.8233 for AUD /USD.

For calculating forward rate for Euro/AUD, we have to take inverse ratios for (AUD/USD) and then multiply with Euro/USD:

The calculations are as follows:

inverse ratios for AUD/USD: (1/1.8233) / (1/1.7725)

=0.5485/0.5642 USD/AUD

multiply with Euro/USD

=(1.0807*0.5485) / (1.1015*0.5642)

=0.5928/0.6215

Thus the 3 months forward rate for Euro/AUD is 0.5928/0.6215

  

The swap points are the differences in interest rates ie., the differences between spot rates and forward rates.

The spot rate of Euro/AUD shall be computed same as taking inverse ratio for AUD/USD.

inverse ratios for AUD/USD (Spot): (1/0.7233) / (1/0.7225)

=1.3826/1.3841

multiply with Euro/USD:

=(1.0607*1.3826) / (1.0615*1.3841)

Thus the spot rate for Euro/AUD is 1.4665/1.4692

Hence the swap points for Euro/AUD = (1.4665 - 0.5928) / (1.4692 - 0.6215)

=0.8737/0.8477.

Summary

The 3 months forward rate for Euro/AUD is 0.5928/0.6215.

The swap points for Euro/AUD is 0.8737/0.8477.

(2.)

3 months & 6 months forward rates for JPY/PKR.

Analyzing for USD/JPY 3 months forward:

spot 122.55/122.72

add 3 months premium: 0.10/0.15

3 months forward for USD/JPY is 122.65/122.87 ----------------{A}

Analyzing for USD/JPY 6 months forward:

spot 122.55/122.72

add 6 months premium: 0.15/0.20

6 months forward for USD/JPY is 122.70/122.92.---------------{B}

Analyzing for USD/PKR 3 months forward:

spot 106.10/106.20

add 3 months premium: 5.90/6.00

3 months forward for USD/PKR is 112.00/112.20---------------{C}

Analyzing for USD/PKR 6 months forward:

spot   106.10/106.20

add 6 months premium: 6.10/6.20

6 months forward for USD/PKR is 112.20/112.40---------------{D}.

To calculate JPY/PKR, we have to inverse USD/JPY and the multiply with USD/PKR.

Calculations are as follows:

3 months forward rate for JPY/PKR = 1/{A} * {C}.

=(1/122.87*112.00) / (1/122.65*112.20)

=0.9115/0.9148

Similarly, 6 months forward rate for JPY/PKR = 1/{B} * {D}

=(1/122.92*112.20) / (1/122.70*112.40)

=0.9128/0.9161

Conclusions:

3 months forward rate for JPY/PKR =0.9115/0.9148

6 months forward rate for JPY/PKR =0.9128/0.9161.

Please note that while taking inverse ratios, the rates have to be interchanged since the "ASK rate" should be greater than the " BID rate".

**Please give me thumps if you find this anwer helpful.